Nine cards numbered 1 to 9 are in a box. A card is drawn and replaced. Then, a second card is drawn. What is the probability that one is prime and the other is even?
Answer by stanbon(75887)

Nine cards numbered 1 to 9 are in a box. A card is drawn and replaced. Then, a second card is drawn. What is the probability that one is prime and the other is even?
----------
Primes: 2,3,5,7
Evens: 2,4,6,8
------
P(one prime and one even) = (4/9)(4/9) = 16/81
==================================================
Cheers,
Stan H.
==============